# Beyond Flat Text: Dual Self-inherited Guidance for Visual Text Generation

### v1

@@ Abstract
In real-world images, slanted or curved texts, especially those on cans, banners, or badges, appear as frequently, if not more so, than flat texts due to artistic design or layout constraints. While high-quality visual text generation has become available with the advanced generative capabilities of diffusion models, these models often produce distorted text and inharmonious text background when given slanted or curved text layouts due to training data limitation. In this paper, we introduce a new training-free framework, STGen, which accurately generates visual texts in challenging scenarios (e.g., slanted or curved text layouts) while harmonizing them with the text background. Our framework decomposes the visual text generation process into two branches: (i) Semantic Rectification Branch, which leverages the ability in generating flat but accurate visual texts of the model to guide the generation of challenging scenarios. The generated latent of flat text is abundant in accurate semantic information related both to the text itself and its background. By incorporating this, we rectify the semantic information of the texts and harmonize the integration of the text with its background in complex layouts. (ii) Structure Injection Branch, which reinforces the visual text structure during inference. We incorporate the latent information of the glyph image, rich in glyph structure, as a new condition to further strengthen the text structure. To enhance image harmony, we also apply an effective combination method to merge the priors, providing a solid foundation for generation. Extensive experiments across a variety of visual text layouts demonstrate that our framework achieves superior accuracy and outstanding quality.
@@ 1 Introduction
Visual text generation is an emerging yet challenging research area in image generation, because text is fine-grained and difficult to be balanced with image. There are two primary methodologies for visual text generation: local visual text blending and global visual text generation. Local visual text blending methods solely render visual text within given images, represented by early GAN-based methods [3, 33, 22, 12, 31]. Recent SceneVTG [38] can render texts in small regions by utilizing diffusion-based models. In contrast, global visual text generation creates new images with integrated text, raising an inevitable challenge to achieve both accurate visual texts and harmonious image content. Recent studies [5, 28, 1, 15, 18, 34, 16, 32] have attempted to address this challenge by leveraging the capabilities of diffusion models [9, 25, 30, 6, 20, 28, 26, 27, 4]. Among them, AnyText [32] distinguishes itself by producing impressive images integrated with outstanding multilingual text. It opens an era of universal visual text generation using large pretrained Visual Text Generation Model (VTGM). In this paper, we follow in the footsteps of AnyText and focus on global visual text generation.
Although large pretrained VGTMs take a significant step towards universal visual text generation, they still do not possess the capability that is robust to handle variable user input, because of insufficient training data. Their performance is unsatisfactory in challenging cases, such as commonly seen slanted or curved texts in real-world images due to artistic design or layout constraints. Specifically, the subpar performance is reflected in text distortion and background occlusion as shown in Fig. 2.
One naive solution is to train VTGM on a more diversified dataset that covers various text configurations. But it is resource-intensive and time-consuming. To overcome this shortcoming, we propose to use VTGM which is well-trained on simple visual text (e.g. flat visual texts) to guide difficult visual text generation (e.g. slanted or curved visual texts). When plainly using the pretrained VTGM (e.g. AnyText), it can generate a latent that contains clear and readable visual texts even at an early stage given a flat text mask. Nevertheless, in tilted situations, text in the latent becomes blurry and wrong, leading to incorrect results.
@@
Figure 2: Failure cases of AnyText [32]. In the top row, we show examples of 2 failure scenarios: Textual distortion (left) and Background occlusion (right). In the bottom row, we show images obtained when using our method under the same seeds.
@@
To address this issue, we propose a training-free method named Slanted Text Generation (STGen). STGen is a plug-and-play method that corrects text regions in latent space during inference. Specifically, our approach employs a dual-branch framework. The first branch, the Semantic Rectification Branch (SRB), utilizes a latent generated using the same prompt, but with a simplified shape, as a robust semantic prior. This branch simultaneously rectifies distorted text predictions and harmonizes the text with its background. For complex visual text generation, such as text layouts composed of multiple tilted or curved sections, we propose a Divide and Conquer strategy to efficiently reconfigure the text shape and obtain a reasonable semantic prior. The second branch, Structure Injection Branch (SIB), extracts rich structural information from glyphs and injects it into the latent space as a structural prior, further enhancing the accuracy of the visual text. Rather than simply merging the two priors, we adopt a novel combination method for optimized integration. Together, the dual-branch framework offers effective guidance for the generation process. Our method addresses the shortcomings of the existing VTGM, ensuring the generated visual texts are both accurate and well-balanced with the overall image as shown in Fig. 1.
@@
To the best of our knowledge, our method is the first specifically for generating visual texts in complex layouts and achieves the state-of-the-art results without additional training. We conduct thorough experiments to demonstrate the superiority of our method and the effectiveness of each component in generating complex visual texts while enhancing overall image quality and coherence. The main contributions of our work are summarized as follows:
@@ 2 Related Works
Denoising diffusion probabilistic models (DDPMs) [9, 25, 30, 6, 20, 28, 26, 27, 4] have made significant contribution to the field of text-to-image synthesis, enabling the generation of high-quality images. Despite advancements in text-to-image synthesis models, achieving precise control over image details remains challenging, particularly in visual text generation.
@@ 2.1 Local Visual Text Blending
Local Visual Text Blending focuses on producing localized images that contain a single line of text. Methods in this area often require a background or original image for text generation or editing. SynText [7] detects the region in the image that can put texts on it and generate texts in the corresponding area. Following their work, SceneVTG [38] introduces a model to predict the text generating area and specifically designs a framework that takes line masks and word masks separately to help generate texts in different shapes. SceneVTG is less user-friendly because it requires an erased image and lacks support for multilingual text generation. AnyTrans [21] provides an end-to-end method for local visual text blending, it identifies the texts in the image, erases them from the image, and generates the translated texts in the same area. The above methods require an input image to render or generate, which is good for editing, but cannot generate the whole image from scratch.
@@ 2.2 Global Visual Text Generation
Recent studies such as Imagen [28] and e-Diff-I [1] find that simply by replacing the text encoder from CLIP [23] to more capable models like T5 [24] can improve visual text generation. Liu et al. [15] further replace the character-blind text encoder with a character-aware text encoder. Even though replacing text encoders seems to be a straightforward approach, it still struggles to generate complex characters such as Korean, Japanese, and Chinese. GlyphByT5 series [16, 17] employ character-aware ByT5 encoder [14] and a new cross attention mechanism to compute the text region and image region separately. GlyphDraw [18] focuses on generating complex characters such as Chinese and proposes a sound baseline by introducing glyph condition for visual text generation. It uses the glyph image of the text and a location mask as conditions, allowing the model to control both the content and position of the generated text. However, it can generate only one line of text per image. TextDiffuser [5] trains a dedicated layout generation module to specifically generate character-level mask as the condition for the diffusion model, improving performance for Latin texts. But it still cannot generate non-Latin texts. GlyphControl [34] employs ControlNet [36] for visual text generation and uses a rendered glyph image as a control condition, making it possible to enable models like Stable Diffusion to generate texts and maintain their text-to-image capability at the same time. Nonetheless, GlyphControl can only generate texts in straight lines and often produces unwanted texts outside the text region. Following the above works, AnyText [32] proposes a robust and unified framework for visual text generation. It can generate high-quality multilingual texts in images, thanks to the text perceptual loss which takes the correctness of the generated text into account during training. However, AnyText performs poorly for generated text leaning more than $45$ degrees. TextGen [35] analyzes how the control information influences the generation in different timesteps. Diff-Text [37] uses off-the-shelf diffusion model and canny controlnet to generate visual texts, where the canny controlnet utilizes the canny image of text glyph to control visual text generation. Nonetheless, Diff-Text binds text position to predefined objects like ’sign’ or ’billboard’ in the prompt. It may fail if a prompt does not include predefined objects, leading to unnatural images.
@@ 3 Method
Figure 3: Pipeline of our method. Given user input on the leftmost side, which contains a prompt and a mask $l_{p}$ specifying positions for generating visual texts, we first split the $l_{p}$ using Divide and Conquer Strategy and obtain glyph image $l_{g}$ and flat position mask $\tilde{l}_{p}$. Then $\tilde{l}_{p}$ and $l_{g}$ are input to Semantic Rectification Branch and Structure Injection Branch respectively. In Semantic Rectification Branch, based on $\tilde{l}_{p}$, we render flat glyph $\tilde{l}_{g}$ and use them along with prompt and random noise $z_{T}$ to generate the latent with flat visual text. This latent serves as a semantic prior, providing rich semantic information for both the generation of the text and its background. $l_{g}$, on the other hand, is converted into the latent space as a structural prior for structural refinement of the text. Finally, the two prior combined together to guide the generation of the visual text in $l_{p}$.
Given a prompt $y$ describing the image and a challenging position mask $l_{p}$, our goal is to generate an image based on $y$ that incorporates visual text at the specified positions dictated by $l_{p}$. Current visual text generation models struggle to generate visual texts in more challenging $l_{p}$ settings, such as tilted or curved texts layouts.
@@ 3.1 Semantic Rectification Branch
In the initial stages of the DDIM denoising process in the VGTM, clear text is created when applying a flat position mask, while blurry and unclear text is generated when using a slanted mask. The reason for the distortion lies in the semantic drift in the text region when tilted. Motivated by this observation, we inherit the high quality flat text generation capabilities of the existing visual text model, adopting the latent of flat text as a constant reference to rectify the semantic information for challenging text generation.
@@
where $l_{p}$ is the position mask input by user. This operation effectively rectifies the visual text using the accurate text in the $z_{f}$. Thanks to the faithful background semantic information embedded in the $z_{f}$, we significantly reduce the erroneous non-textual semantic information in the text region while maintaining a coherent background, thereby avoiding background occlusion.
@@
AdaIN [10] is originally developed for style transfer tasks. It substitutes mean and standard deviation of the source feature with those of the target feature. Masui et al. [19] demonstrated that AdaIN can be applied directly to Diffusion models for style transfer without any additional training. In our method, we also incorporate AdaIN, but with the goal of preserving image integrity against disruptions in latent distribution caused by the replacement operation:
@@
1 Input: A text prompt $y$, an input position image $l_{p}$, a flat reference position image $\tilde{l}_{p}$ and a set of iterations for refinement $\left\{t_{1},\ldots,t_{k}\right\}$ and a trained Visual Text Generation Model $\mathcal{M}$
2 Output: Denoised vector $z_{0}$
3   Render the glyph image $l_{g}$ and compute its latent $z_{g}$;
4 Compute the reference latent $z^{f}_{0}$;
5 Rotate and crop $z^{f}_{0}$ to get $z_{f}$ according to the transformation from $\tilde{l}_{p}$ to $l_{p}$;
6 for $t=T,T-1,\ldots,1$ do
7       if $t\in\left\{t_{1},\ldots,t_{k}\right\}$ then
8             $\tilde{z}_{t}\leftarrow{\rm AdaIN}(z_{f},z_{t})\odot l_{p}+z_{t}\odot(1-l_{p})$;
9             $\hat{z}_{t}\leftarrow\rho{\rm AdaIN}(z_{g},z_{t})+(1-\rho)\tilde{z}_{t}$;
10             $\ddot{z}_{t}\leftarrow(\kappa_{t}\lambda\hat{z}_{t}+(1-\kappa_{t})z_{t})\odot l_{p}+z_{t}\odot(1-l_{p})$;
11             Set $z_{t}\leftarrow\ddot{z}_{t}$;
12
13       end if
14      Set $z_{t-1}\leftarrow\mathcal{M}(z_{t},y,t,l_{p},l_{g})$;
15
16 end for
17
18Return $z_{0}$
Algorithm 1 STGen
@@ 3.2 Structure Injection Branch
where $\lambda$ is a hyper-parameter and the $\kappa_{t}$ is a temporal factor that decays over time with each timestep. Together, they control the injection strength of the merged prior. The complete algorithm is represented as Algorithm 1.
@@ 4 Experiments
Table 2: Quantitative Comparison between STGen and other competitors on both English and Chinese sets. All competitors are evaluated based on their officially released models. Numbers in bold indicate the best performance, and underscored numbers indicate the second best.
Baseline Baseline Preference Our Method Preference
TextDiffuser 23.07% 76.93%
GlyphControl 21.96% 78.04%
SD1.5+SceneVTG 18.99% 81.01%
AnyText 25.31% 74.69%
Diff-Text 17.00% 83.00%
Table 3: User study results. The participants were asked to select the best results based on image quality, the accuracy of the generated visual texts, and the similarity between the prompt and the generated images.
@@ 4.1 Implementation Details
Our framework is based on the pretrained AnyText model. During inference, we utilize a single RTX 3090 GPU for images in size of (512, 512). The framework achieves optimal performance when $\lambda$ is within the range $[-0.5,0.5]$ and $\rho$ is within $(0,2]$. We set $\kappa_{t}=10^{t-T}$ for effective guidance and harmonious text region boundary. Specifically, when $\lambda$ is in $[-0.5,0)$, the diversity of generated visual texts increases, while $\lambda$ in $(0,0.5]$ increases the variety of the image background. As shown in Tab. 1, our method demonstrates robustness with respect to hyper-parameter sensitivity. In subsequent experiments, we set $\lambda$ to $0.5$ and $\rho$ to $0.5$ which has a balanced performance on both Chinese and English for evaluation.
@@ 4.2 Evaluation Setup
Since there are currently no openly available datasets that focus on challenging visual texts generation, we establish a new benchmark for evaluation based on the AnyText-benchmark [32]. For each text position mask, we randomly rotate the masks from the original benchmark and resolve any intersections, resulting in 984 prompts for LAION-word (English evaluation) and 919 prompts for Wukong-word (Chinese evaluation), each with corresponding position masks. Each prompt includes at least one visual text line and up to five visual text lines. During evaluation, we categorize the masks into three levels based on their angles: ‘easy’ (0°–30°), ‘medium’ (30°–60°), and ‘hard’ (60°–90°), facilitating multi-level assessments.
@@
We evaluated existing competing methods, including TextDiffuser [5], GlyphControl [34], SceneVTG [38], AnyText [32] and Diff-Text [37] using the benchmark and metrics mentioned above. It is worth noticing that the SceneVTG [38] cannot generate image from scratch. Therefore, we use the images generated by Stable Diffusion [27] with the same prompt as background image. For all the methods in the evaluation, we use the preprocessed mask using the Divide and Conquer strategy mentioned above to generate images to ensure fairness.
Figure 4: Qualitative comparison of our method and state of the art models in both English and Chinese text generation.
4.3 Qualitative Comparisons
As presented in Fig. 4, TextDiffuser tends to generate low-quality images, specifically in the second and the fifth columns. GlyphControl often produces texts outside the given text position, as can be seen in the fourth column. SceneVTG [38] exhibits severe text distortion, particularly in the first and third columns where the texts are extremely distorted and unrecognizable. Typical background occlusion is also observed in the second, third, and fourth columns, where the texts “Burning”, “Mamba” and “Twinkle” are obscured by non-textual elements such as lava, a human figure, and stars, respectively. This indicates that SceneVTG can only render texts against simple and clear backgrounds and does not integrate well with other text-to-image models. For Diff-text [37], there is a noticeable lack of coherence in the images due to insufficient consideration for harmony between the visual texts and the background, as particularly evidenced in the last 2 columns where the texts are rendered outside the bow and the watch. In AnyText [32], the generated visual texts tend to be distorted, this was caused by the loss of semantic and structure information during the inference. Moreover, in the first column, the text “Track” occludes with the train head, affecting both the text and the train. For our method, we can generate images with superior quality with accurate English and Chinese visual texts integrated even in complex scenarios like in the fourth and fifth columns. It’s worth noting that except for resolving text distortion, our method also improves the background occlusion between visual text and non-textual elements. For instance, in the third column of Fig. 4, the text “Mamba” is originally disrupted by the human head using AnyText [32], while our method pushes the human head aside to generate a clear text and reasonable non-textual element.

### v2

@@ Abstract
In real-world images, slanted or curved texts, especially those on cans, banners, or badges, appear as frequently, if not more so, than flat texts due to artistic design or layout constraints. While high-quality visual text generation has become available with the advanced generative capabilities of diffusion models, these models often produce distorted text and inharmonious text backgrounds when given slanted or curved text layouts due to training data limitations. In this paper, we introduce a new framework, STGen, which accurately generates visual texts in challenging scenarios (e.g., slanted or curved text layouts) while harmonizing them with the text background. Our framework decomposes the visual text generation process into two branches: (i) Semantic Rectification Branch, which leverages the ability in generating flat but accurate visual texts of the model to guide the generation of challenging scenarios. The generated latent of flat text is abundant in accurate semantic information related both to the text itself and its background. By incorporating this, we rectify the semantic information of the texts and harmonize the integration of the text with its background in complex layouts. (ii) Structure Injection Branch, which reinforces the visual text structure during inference. We incorporate the latent information of the glyph image, rich in glyph structure, as a new condition to further strengthen the text structure. To enhance image harmony, we also apply an effective combination method to merge the priors, providing a solid foundation for generation. Extensive experiments across a variety of visual text layouts demonstrate that our framework achieves superior accuracy and outstanding quality. The code will be available at https://github.com/Tony-Lowe/STGen.
@@ 1 Introduction
Visual text generation is an emerging yet challenging research area in image generation because text is fine-grained and difficult to balance with the image. The current methods [3, 37, 25, 15, 35, 43, 6, 31, 1, 18, 21, 38, 19, 36, 5] proposed a series techniques to address it. Among them, diffusion based methods (AnyText [36], GlyphControl [38], DiffText [41] etc.) can create new images with integrated text, raising an inevitable challenge to achieve both accurate visual texts and harmonious image content. Notably, AnyText [36] distinguishes itself by producing impressive images integrated with outstanding multilingual text. It opens an era of universal visual text generation using large pretrained Visual Text Generation Model (VTGM).
Although large pretrained VGTMs take a significant step towards universal visual text generation, they still struggle to handle users’ diverse inputs, such as commonly seen slanted or curved texts in real-world images. Given these user inputs, the model often leads to text distortion and background occlusion, as shown in Fig. 2. This is because the latent space gradually becomes blurry and distorted (as shown in Fig. 3) due to insufficient data in such scenarios, thus the model cannot effectively maintain the structure and semantic information in the text region as it does when processing a flat mask.
One naive solution is to train VTGM on a more diversified dataset that covers various text configurations. But it is resource-intensive and data distribution across multiple text layouts cannot be guaranteed. Even if they are trained on those data, the results may not be satisfying [32, 10].
@@
Figure 2: Failure cases of AnyText [36]. The top row illustrates two failure cases: textual distortion (left) and background occlusion (right). The bottom row displays results using our method.
@@
To overcome this shortcoming, we propose a plug-and-play method named Slanted Text Generation (STGen), which corrects text regions in latent space during inference. Specifically, our approach employs a dual-branch framework. The first branch, the Semantic Rectification Branch (SRB), utilizes a latent generated using the same prompt, but with a simplified shape, as a robust semantic prior. This branch simultaneously rectifies distorted text predictions and harmonizes the text with its background. For complex visual text generation, such as text layouts composed of multiple tilted or curved sections, we propose a Divide and Conquer strategy to efficiently reconfigure the text shape and obtain a reasonable semantic prior.
The second branch, Structure Injection Branch (SIB), extracts rich structural information from glyphs and injects it into the latent space as a structural prior, further enhancing the accuracy of the visual text. Rather than simply merging the two priors, we adopt a novel combination method for optimized integration for better coherence in latent space. Together, the dual-branch framework offers effective guidance for the generation process without requiring additional training.
@@
To the best of our knowledge, our method is the first tailored specifically for generating visual texts in complex layouts and achieves state-of-the-art results as shown in Figs. 1, 5 and 6 without additional training. Our framework effectively improves visual text accuracy across various approaches, including early-stage GlyphControl, DiffText, which receives no specified training for visual text generation, and AnyText. We conduct thorough experiments to demonstrate the superiority of our method and the effectiveness of each component in generating complex visual texts while enhancing overall image quality and coherence. The main contributions of our work are summarized as follows:
@@
Figure 3: Comparison of predicted $x_{0}$ under different inference steps. The first and second rows show AnyText’s intermediate predicted $x_{0}$ for flat and slanted masks, respectively. While predictions remain stable under flat masks, the $x_{0}$ prediction drifts during inference with slanted masks. Our method effectively guides the model to maintain accuracy in visual texts.
@@ 2 Related Works
Denoising diffusion probabilistic models (DDPMs) [12, 28, 34, 8, 23, 31, 29, 30, 4] have made significant contribution to the field of text-to-image synthesis. However, achieving precise control over image details remains challenging, particularly in visual text generation.
@@ 2.1 Local Visual Text Blending
Local Visual Text Blending methods synthesize localized images with single-line textual content. Existing approaches typically necessitate a background image for text insertion or modification. SynText [9] identifies image regions suitable for text placement and renders text accordingly. Following their work, SceneVTG [43] introduces a model to predict the text-generating area and specifically designs a framework that takes line masks and word masks separately to help generate texts in different shapes. SceneVTG is less user-friendly because it requires an erased image and lacks support for multilingual text generation. AnyTrans [24] detects, removes, and replaces text with translations in the same region. The above methods require an input image to render or generate, which is good for editing, but cannot generate the whole image from scratch.
@@ 2.2 Global Visual Text Generation
Recent studies, such as Imagen [31], demonstrate that replacing CLIP text encoder [26] with more advanced models like T5 [27] enhances visual text generation. Liu et al. [18] further replace the character-blind text encoder with a character-aware text encoder. GlyphByT5 series [19, 20] employ character-aware ByT5 encoder [17] and a new cross attention mechanism to compute the text region and image region separately. Although replacing text encoders appears straightforward, it still struggles to generate complex characters such as Korean, Japanese, and Chinese. GlyphDraw [21] pioneered using glyph conditions and location masks for visual text generation of complex characters. It employs a glyph image and a location mask to control content and placement, though it remains limited to one line per image. TextDiffuser [6] introduces a dedicated layout generation module to produce character-level masks as diffusion model conditions, enhancing Latin text generation. But it still cannot generate non-Latin texts. TextDiffuser-2 [5] refines this approach by replacing character-level masks with bounding box coordinates, but this modification weakens its capacity for flexible visual text layout customization. GlyphControl [38] leverages ControlNet [40] and uses rendered glyph images as control conditions, enabling visual text generation in models like Stable Diffusion while preserving their core image synthesis capabilities. Nevertheless, GlyphControl is restricted to straight-line text layouts and frequently generates extraneous text artifacts. Building on these works, AnyText [36] proposes a unified framework for high-quality multilingual visual text generation. Its text perceptual loss directly evaluates text accuracy during training, improving correctness of generated visual text. However, AnyText struggles with rendering text rotated beyond 45 degrees. TextGen [39] examines how control signals affect image generation across timesteps. TextHarmony [42] unifies image comprehension, generation, and editing within a single framework by training a multimodal Large Language Model (LLM) to improve image understanding and generate more precise tokens for image synthesis. However, users face challenges in defining text layout during image generation, as current methods rely solely on textual prompts. In terms of text editing, its performance in complex backgrounds is unsatisfactory. Diff-Text leverages Canny ControlNet for glyph-based generation but ties text placement to predefined objects in prompts. It may fail if a prompt does not include predefined objects, leading to unnatural images.
@@ 3 Method
Figure 4: Pipeline of our method. Given user input on the leftmost side, which contains a prompt and a mask $l_{p}$ specifying positions for generating visual texts, we first split the $l_{p}$ using Divide and Conquer Strategy and obtain glyph image $l_{g}$ and flat position mask $\tilde{l}_{p}$. Then $\tilde{l}_{p}$ and $l_{g}$ are input to the Semantic Rectification Branch and Structure Injection Branch respectively. In the Semantic Rectification Branch, based on $\tilde{l}_{p}$, we render flat glyph $\tilde{l}_{g}$ and use them along with prompt and random noise $z_{T}$ to generate the latent with flat visual text. This latent serves as a semantic prior, providing rich semantic information for both the generation of the text and its background. $l_{g}$, on the other hand, is converted into the latent space as a structural prior for structural refinement of the text. Finally, the two prior combined to guide the generation of the visual text in $l_{p}$.
Given a prompt $y$ describing the image and a challenging position mask $l_{p}$, our goal is to generate an image based on $y$ that incorporates visual text at the specified positions dictated by $l_{p}$. Current visual text generation models struggle to generate visual texts in more challenging $l_{p}$ settings, such as tilted or curved text layouts.
@@ 3.1 Semantic Rectification Branch
In the initial stages of the DDIM denoising process in the VGTM, clear text is generated. This text remains accurate when a flat position mask is applied. However, when using a slanted mask, the text gradually becomes distorted. The reason for the distortion lies in the semantic drift in the text region when tilted. Motivated by this observation, we inherit the high-quality flat text generation capabilities of the existing visual text model, adopting the latent of flat text as a constant reference to rectify the semantic information for challenging text generation.
@@
where $l_{p}$ is the position mask input by the user. This operation effectively rectifies the visual text using the accurate text in the $z_{f}$. Thanks to the faithful background semantic information embedded in the $z_{f}$, we significantly reduce the erroneous non-textual semantic information in the text region while maintaining a coherent background, thereby avoiding background occlusion.
@@
AdaIN [13] is originally developed for style transfer tasks. It substitutes mean and standard deviation of the source feature with those of the target feature. Masui et al. [22] demonstrated that AdaIN can be applied directly to Diffusion models for style transfer without any additional training. In our method, we also incorporate AdaIN, but to preserve image integrity against disruptions in latent distribution caused by the replacement operation:
@@ 3.2 Structure Injection Branch
where $\lambda$ is a hyper-parameter and the $\kappa_{t}$ is a temporal factor that decays over time with each timestep. Together, they control the injection strength of the merged prior.
@@ 4 Experiments
Table 2: Quantitative Comparison between STGen and other competitors on both English and Chinese sets. All competitors are evaluated based on their officially released code and models. Numbers in bold indicate the best performance, and underscored numbers indicate the second best. indicates we adapt the method which is originally focused on other tasks into image generation.
Method SEN.ACC $\uparrow$ NED $\uparrow$ FID $\downarrow$
GlyphControl 37.10/3.27 66.80/8.45 37.84/34.36
w/ STGen 70.92/11.81 84.68/24.72 32.00/32.99
Diff-Text 56.11/29.49 75.28/48.67 69.24/62.61
w/ STGen 68.55/39.58 81.55/55.61 66.77/61.59
AnyText 72.39/69.23 87.67/83.96 33.54/31.58
w/ STGen 75.53/70.00 89.38/84.57 33.83/32.17
Table 3: Evaluation on vanilla AnyText Benchmark. In each cell of the table, the numbers on the right are results from the English set while the numbers on the right are from the Chinese set.
@@ 4.1 Implementation Details
We use a single RTX 3090 GPU for images of size (512, 512). Optimal performance is achieved when $\lambda\in[-0.5,0.5]$ and $\rho\in(0,2]$. We set $\kappa_{t}=10^{t-T}$ for effective guidance and harmonious text region boundary.As shown in Tab. 1, our method demonstrates robustness for hyper-parameter sensitivity. In subsequent experiments, we set $\lambda$ to $0.5$ and $\rho$ to $0.5$, which has a balanced performance on both Chinese and English for evaluation.
@@ 4.2 Evaluation Setup
Due to the lack of publicly available datasets focused on challenging visual text generation, we propose a new benchmark derived from the AnyText-benchmark [36]. For each text position mask, we randomly rotate the original benchmark’s masks and resolve overlaps, yielding 984 prompts for LAION-word (English evaluation) and 919 prompts for Wukong-word (Chinese evaluation). During evaluation, masks are categorized into three difficulty levels based on rotation angles: easy (0°–30°), medium (30°–60°), and hard (60°–90°), enabling multi-level assessment of performance.
@@
We evaluated existing competing methods, including TextDiffuser [6], TextDiffuser-2 [5], TextHarmony [42], GlyphControl [38], SceneVTG [43], AnyText [36] and Diff-Text [41] using the benchmark and metrics mentioned above. Notably, SceneVTG [43] cannot generate images from scratch. To address this, we use background images produced by Stable Diffusion [30] with identical prompts. For TextHarmony [42], its image generation model lacks support for positioning visual text at specific locations. Following its evaluation protocol on the AnyText Benchmark, we employ its visual text editing mode, which generates text on images where target positions are masked in black. For all the baselines in the evaluation, we use their officially released code and checkpoints. To ensure fairness, we preprocess the masks using the Divide and Conquer strategy mentioned above before generating images.
Baseline Baseline Preference Ours Preference
TextDiffuser [6] 23.07% 76.93%
TextDiffuser-2 [5] 20.15% 79.85%
GlyphControl [38] 21.96% 78.04%
SceneVTG [43] 18.99% 81.01%
TextHarmony [42] 18.75% 81.25%
AnyText [36] 25.31% 74.69%
Diff-Text [41] 17.00% 83.00%
Table 4: User study results. Participants were asked to choose the best results based on image quality, accuracy of generated text within images, and prompt-image similarity.
Figure 5: Qualitative comparison of our method and state-of-the-art models in both English and Chinese text generation.